8. Newton’s rings formed by monochromatic light between a flat glass plate and a plano-convex lens are viewed normally. Calculate the order of the dark ring which will have double the diameter of that of 40th dark ring.
1
Expert's answer
2019-08-26T10:40:59-0400

For the given configuration, the radius of the nth dark ring is given by:


"r_n^2=nR\lambda"

So, the diameter is given by:


"d_n=\sqrt{4nR\lambda}"

So, the diameter twice as the diameter of the 40th dark ring has:


"n=n_{40}*(\frac{d}{d_{40}})^2=40*2^2=40*4=160"

So, the 160th order dark ring has the diameter twice as big as diameter of the 40th dark ring